OpenStudy (anonymous):

What is 4.6 × 108 cm/min in meters per hour? A. 2.76 × 108 m/h B. 2.76 × 1010 m/h C. 7.67 × 106 m/h D. 7.67 × 108 m/h

OpenStudy (anonymous):

4.6 × 10^8 cm/min = 4.6 × 10^8 x 60 / 100 meters per hour = 2.76 x 10^8 meters pe hour. So the answer is A

OpenStudy (anonymous):

oh thank you